[Brahms-l] Shift report 20050424 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Apr 24 2005 - 23:52:16 EDT
Shifter leader was kh
Shifters were 


Machine performed well.  Beam was dumped at 19:00, we had an access to 
replace LR4032a and beam was back at 21:00.  The changeout of the LR4032 was successful this time and hence we are back to running with the full complement of detectors.
Runs achieved
Run   tr2   tr3
No CCR1-5, BBR1-15, or MRST0
14244  42483  off
14245  40943  off
14246  34474  off
LR4032 replaced; CC1-5 and other detectors on after this
14250  52447   83314
14251  56596  175741
14252 in progress

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Apr 24 23:52:29 2005

This archive was generated by hypermail 2.1.8 : Sun Apr 24 2005 - 23:52:50 EDT